OpenStudy (johnnewman):

A six-sided number cube with faces numbered 1, 2, 3, 4, 5, and 6 is rolled. What is the probability of rolling a 2 or a number greater than 5? A. 1/36 B. 1/3 C. 1/2 D. 2/3 <---

OpenStudy (johnnewman):

Its D

OpenStudy (smartnerd1111):

No its b becuasss you have 2 and you have 6 citing in the description so 2 numbers out of 6 is also equal to 1/3

OpenStudy (smartnerd1111):

Please meal me

OpenStudy (smartnerd1111):

Medal is what I meant

OpenStudy (mrnood):

D is not correct

OpenStudy (trojanpoem):

rolling a 2 1/6 a number > 5 6 1/6 so you have P(A) = 1/6 + 1/6 = 2/6 = 1/3

OpenStudy (phan_trash):

Its actual;;y B

OpenStudy (phan_trash):

*actually